OpenStudy (jazzyfa30):

Which equation has a graph that is perpendicular to the graph of 4x - 2y = 1? a. y = -2x + 8 b. 2x + 4y = -1 c. y = 2/3x - 2 d. 6x -3y = 9

OpenStudy (phi):

First, use algebra to change 4x - 2y = 1 into y = mx + b form can you do that ?

OpenStudy (jazzyfa30):

-2y=-4x+1

OpenStudy (phi):

yes, so far, so good, now divide both sides and all terms by -2 what do you get ?

OpenStudy (jazzyfa30):

y=2x+1/-2

OpenStudy (jazzyfa30):

I did divide already 1/-2is a fraction

OpenStudy (phi):

no , you are correct. But I would write it as y = 2x - 1/2

OpenStudy (jazzyfa30):

oh ok

OpenStudy (phi):

to find the slope of a line perpendicular to y = 2x - 1/2 you take the slope (the 2 which is in front of the x) multiply by -1 and then "flip it" to get -1/2 that is the slope you want in the answer.

OpenStudy (phi):

the slope is m in y= mx + b in choice (a), the slope is -2. So that is not the answer neither is (c)

OpenStudy (phi):

to see if (b) is the answer, put 2x + 4y = -1 into y= mx+b form

OpenStudy (jazzyfa30):

4y=-2x-1 (/4) y=-2x-1/4

OpenStudy (phi):

and the slope is ?

OpenStudy (jazzyfa30):

-1/2 so my answer is b

OpenStudy (phi):

yes

OpenStudy (jazzyfa30):

thank you for both the medal and the help and here the next one Which equation has a graph that is parallel to the graph of 4x - 2y = 1? Select one: a. y = 2/3x - 2 b. y = -2x + 8 c. 2x -+ 4y = -1 d. 6x -3y = 9

OpenStudy (phi):

parallel means the 2 lines have the same slope. so put 4x - 2y = 1 into y= mx+b form

OpenStudy (jazzyfa30):

-2y=-4x+1

OpenStudy (phi):

now divide both sides by -2

OpenStudy (jazzyfa30):

y=2x-1/2

OpenStudy (phi):

and the slope is ?

OpenStudy (jazzyfa30):

1/2

OpenStudy (phi):

no the slope is m in y= mx +b match "m" to the number in front of the x

OpenStudy (jazzyfa30):

oh 2

OpenStudy (phi):

now you have to find the slope of each line in the choices. and find the one that has a slope of 2

OpenStudy (jazzyfa30):

I think its between b and d

OpenStudy (phi):

in (b) you have y = -2x + 8 the number in front of the x is -2. so the slope is -2. that is not the same as 2

OpenStudy (jazzyfa30):

so its between c & d

OpenStudy (phi):

so put each into y = mx + b form and see which one has m=2 (the slope = 2)

OpenStudy (jazzyfa30):

the answer's d